OpenStudy (anonymous):

Solve the system of equations by elimination: 12x-10y=0 -6x+5y=2

OpenStudy (anonymous):

again, there is no solution

OpenStudy (anonymous):

Why? or How

OpenStudy (anonymous):

u end up with 4=0

OpenStudy (anonymous):

notice the slope. they are the same

OpenStudy (anonymous):

Will you show me the work please?

OpenStudy (anonymous):

12x-10y =0 can be re rewritten as (dividing by 2 on both sides) 6x-5y = 0 (multiply by -1 on both sides) or -6x+5y = 0

OpenStudy (anonymous):

so your two equations are -6x+5y =0 ------>1 -6x+5y =2 ------->2 subtract 2 from 1 0 =-2 this is nonsensical so this has no solution.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

What about 4x-6y=20 2x-3y=10

OpenStudy (anonymous):

this = 0?

OpenStudy (anonymous):

they are the same line. they have infinitely many solutions

OpenStudy (anonymous):

So? what do i write

OpenStudy (anonymous):

infinitely many solutions

jhonyy9 (jhonyy9):

6x-5y=0 --- 6x=5y --- x=5y/6 -6x+5y=2 -- (-6(5y/6)) + 5y =2 --- -5y +5y=2 --- 0=2